# Multi-modal Misinformation Detection: Approaches, Challenges and Opportunities

### v6

@@ 5.1. Data related challenges
  • Lack of large and comprehensive datasets: As illustrated in Fig. 4, most of the existing datasets are small in size, and sometimes highly imbalanced in terms of fake to real ratio.
  • Lack of cross-lingual datasets: Almost all social media platforms are multi-lingual environments where users share information in multiple languages. Although misinformation spreads in multiple languages, a vast majority of the existing datasets are mono-lingual i.e., they only provide English content. Therefore, there is a serious lack of non-English content and annotations.
  • Limited modalities: As discussed earlier, most of the existing multi-modal datasets only provide image and text modalities. Thus, neglecting useful information that are conveyed by other modalities such as video, audio etc. The necessity of providing more modalities becomes more apparent when we consider popular social media such as YouTube, TikToK and Clubhouse that are mainly video or audio based platforms.
  • Bias in event-specific datasets: Many of the existing datasets are created for specific events such as COVID-19 crises, thereby not covering a variety of events and topics and as a result of this they may not sufficiently train models to detect fake news in other contexts.
  • Binary and domain level ground truth: Most of the existing datasets provide us with binary and domain level ground truth for well-known outlets such as Onion or The New York Times. In addition, they often do not provide any information about reasons of misinformation e.g., cross-modal discordance, false connection, imposter content etc.
  • Subjective annotations and inconsistency of labels: As discussed in the data study section, different datasets use different crowd-source and fact-checking agencies, thereby articles are annotated subjectively with different labels across different datasets. Thus, it is very challenging to analyze, compare and interpret results.
5.2. Feature related challenges:
This category comprises shortcomings related to the cross-modal feature identification and extraction in the multi-modal fake news detection pipeline. Some of the most important weaknesses in current feature based studies are:
  • Insufficiency of cross-modal cues: Although researchers have proposed some multi-modal cues, most of the existing models naively fuse image based features with textual features as a supplement. There are fewer works that leverage explainable cross-modal cues other than image and text combination. However, there are still plenty of useful multi-modal cues which are often neglected by researchers.
@@ 5.2. Feature related challenges:
  • Lack of language-independent features: The majority of existing work on misinformation leverages text features that are highly dependent on dataset languages which is mostly English. Identifying language-independent features, is an effective way to cope with mono-lingual datasets.
5.3. Model related challenges
@@
  • Inexplicability of current models: A majority of the existing models do not provide any explicable information about the regions of interest, common patterns of inconsistencies among modalities and type of misinformation e.g., manipulation, exaggeration etc. While some recent works attempt to use attention-based techniques to overcome the problem of ineffective multi-modal embedding and provide some interpretability, most of them usually follow a trial and error approach like masking to find relevant sections to attend to. However, interpretable and explainable AI is crucial in building trust and confidence as well as fairness and transparency which are being mostly neglected.
  • Non-transferable models to unseen events: Most of the existing models are designed in such a way that they extract and learn event-specific features e.g., COVID-19, election etc. Thus, they are most likely biased toward specific events and as a result, not transferable to unseen and emerging events. For this very reason, building models that learn general features and separate them from the nontransferable event-specific features would be extremely useful.
  • Unscalability of current models: Considering the expensive and complicated structures of deep networks and the fact that most of the existing multi-modal models leverage multiple deep networks (one for each modality), they are not scalable if the number of modalities increases. Moreover, many of the existing models require heavy computing resources and need large volume of memory storage and processing units. Therefore, scalability of proposed models should be taken into account while developing new architectures.
@@
5.4. Opportunities in multi-modal misinformation detection study
@@
Considering the challenges and shortcomings in multi-modal misinformation detection we discussed above, we propose opportunities in furthering research in this field. In what follows, we discuss these opportunities by each direction of multi-modal misinformation detection study.
5.4.1. Opportunities in multi-modal data study
@@
Considering the data related challenges we discussed earlier, we propose the following data related avenues:
@@
  • Comprehensive multi-modal and multi-lingual datasets: As we discussed in detail, one important gap in the misinformation detection study is the lack of a comprehensive multi-modal dataset which needs to be addressed in the future. Multi-modal misinformation detection requires large, multi-lingual, multi-source datasets that cover a variety of modalities, web resources, events etc. and provide fine-grained ground truth for the samples.
@@
  • Standardized annotation strategy: As mentioned above, current datasets annotate articles by different fact-checking agencies and as a result the labels are in many cases subjective. Having a standard agreement across all datasets on how to label articles makes the cross-dataset comparison and analysis much easier.
5.4.2. Opportunities in multi-modal feature study
@@
: Based on feature related challenges we discussed in the previous section, we propose the following research opportunities to overcome some of the existing challenges in multi-modal feature study:
  • Identifying cross-modal clues: As mentioned earlier, cross-modal cues are currently limited to a handful of trivial clues such as similarity of text and image. Identifying subtler, yet neglected cues not only helps in development of discordance-aware models, but also could be helpful in recognizing vulnerabilities of the serving platforms which is a part and parcel of adversarial learning.
@@
  • Developing efficient fusion mechanism: As discussed before, many of the existing solutions leverage naive fusion mechanisms such as concatenation which may result in inefficient and noisy multi-modal representations. Therefore, another fruitful avenue of research lies in the study and development of more efficient fusion techniques to produce information richer representations.
@@
  • Identifying language-independent features to cope with mono-lingual datasets: As discussed above, the majority of existing datasets are mono-lingual, thereby are not sufficient enough to train models for non-English tasks. One way to compensate for the lack of multi-lingual datasets is to use language-independent features (Vogel and Meghana, 2020). Identifying such features, specially in multi-modal environments where there are more features and aspects, would be highly effective in coping with mono-lingual datasets.
5.4.3. Opportunities in multi-modal model study
@@
In what follows, we propose some unexplored research avenues to tackle some of the existing model related challenges in multi-modal misinformation detection.
@@
  • Utilizing foundation models and prompt-based techniques in multi-modal misinformation detection: The astounding effectiveness of foundation models and techniques, including ICL and prompt-tuning, in numerous multi-modal tasks suggests that foundation models have a lot of potential for identifying multi-modal misinformation. Developing task specific foundation model for detecting misinformation is another opportunity that would hugely impact the field of misinformation detection.
@@
  • Developing cross-modal discordance-aware architectures: As described earlier, most of the existing works, either blindly merge modalities or take a trial and error approach to attend to the relevant modalities. Implementing discordance-aware models not only results in information richer representations, but also may be useful in making attention based techniques more efficient.
@@
  • Adversarial learning in multi-modal misinformation detection: Although there are existing generative-based architectures, adversarial study of multi-modal misinformation detection has been mostly neglected. In order to make the detection models more adversarially robust, it is of utmost importance to dedicate time and effort to the study and development of generative and adversarial learning techniques.
@@
  • Transferable models to unseen events: As mentioned earlier, except a few works, most of the existing models are designed for specific events and as a result, are ineffective for emerging ones. Since misinformation spreads during a variety of events, developing general and transferable models is extremely crucial.
@@ 6. Conclusions
In this paper, we review the literature on multi-modal misinformation detection, discuss its strengths and weaknesses, and suggest new directions for future research. First, we introduce some of the prominent misinformation categories and often used cross-modal cues for spotting them. We also discuss different fusion mechanisms to merge modalities that are engaged in such cross-modal cues. In addition, we categorize existing solutions into two groups: classic machine learning and deep learning solutions, and then further divide each group based on the techniques that are utilized. Furthermore, we introduce and compare existing datasets on multi-modal misinformation detection and dentify some of the weaknesses of these datasets. By classifying them into data, feature, and model-based shortcomings, we demonstrate some of the most prominent problems in multi-modal fake news detection. Finally, we propose new lines of research to address these shortcomings.
7. Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant number 2127309 to the Computing Research Associate for the CIFellows Project. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author and do not necessarily reflect the views of funding agencies.

### v7

@@ 5.1. Data study Challenges
  • Lack of large and comprehensive datasets: as illustrated in Fig. 7, most of the existing datasets are small in size and sometimes highly imbalanced in terms of the fake-to-real ratio.
  • Lack of cross-lingual datasets: almost all social media platforms are multi-lingual environments where users share information in multiple languages. Although misinformation spreads in multiple languages, a vast majority of the existing datasets are mono-lingual, i.e., they only provide English content. Therefore, there is a serious lack of non-English content and annotations.
  • Limited modalities: as we discussed earlier, most of the existing multi-modal datasets only provide image and text modalities, thus neglecting useful information conveyed by other modalities such as video, audio, etc. The necessity of providing more modalities becomes more apparent when we consider popular social media such as YouTube, TikTok, and Clubhouse, which are mainly video or audio-based platforms.
  • Bias in event-specific datasets: many of the existing datasets are created for specific events such as the COVID-19 crisis, thereby not covering a variety of events and topics. As a result, they may not sufficiently train models to detect fake news in other contexts.
  • Binary and domain-level ground truth: most of the existing datasets provide binary and domain-level ground truth for well-known outlets such as The Onion or The New York Times. In addition, they often do not provide any information about the reasons for misinformation, e.g., cross-modal discordance, false connection, imposter content, etc.
  • Subjective annotations and inconsistency of labels: as discussed in the data study section, different datasets use different crowd-sourced and fact-checking agencies, thereby articles are annotated subjectively with different labels across different datasets. Thus, it is very challenging to analyze, compare, and interpret results.
5.2. Feature Study Challenges
This category comprises shortcomings related to cross-modal feature identification and extraction in the multi-modal fake news detection pipeline. Some of the most important weaknesses in current feature-based studies are:
  • Insufficiency of cross-modal cues: although researchers have proposed some multi-modal cues, most of the existing models naively fuse image-based features with textual features as a supplement. There are fewer works that leverage explainable cross-modal cues other than image and text combinations. However, there are still plenty of useful multi-modal cues that are often neglected by researchers.
@@ 5.2. Feature Study Challenges
  • Lack of language-independent features: a majority of existing work on misinformation leverages text features that are highly dependent on dataset languages, which are mostly English. Identifying language-independent features is an effective way to cope with mono-lingual datasets.
5.3. Model Study Challenges
@@
  • Inexplicability of current models: a majority of the existing models do not provide any explicable information about the regions of interest, common patterns of inconsistencies among modalities, and types of misinformation (e.g., manipulation, exaggeration). While some recent works attempt to use attention-based techniques to overcome the problem of ineffective multi-modal embedding and provide some interpretability, most of them usually follow a trial-and-error approach like masking to find relevant sections to attend to. However, interpretable and explainable AI is crucial in building trust and confidence as well as fairness and transparency, which are mostly neglected.
  • Non-transferable models to unseen events: most of the existing models are designed in such a way that they extract and learn event-specific features (e.g., COVID-19, election). Thus, they are most likely biased toward specific events and, as a result, not transferable to unseen and emerging events. For this reason, building models that learn general features and separate them from the non-transferable event-specific features would be extremely useful.
  • Unscalability of current models: considering the expensive and complicated structures of deep networks and the fact that most of the existing multi-modal models leverage multiple deep networks (one for each modality), they are not scalable if the number of modalities increases. Moreover, many of the existing models require heavy computing resources and need a large volume of memory storage and processing units. Therefore, the scalability of proposed models should be taken into account while developing new architectures.
@@
6. Opportunities in Multi-modal Misinformation Detection
@@
Considering the challenges and shortcomings in multi-modal misinformation detection we discussed above, we propose opportunities for furthering research in this field. In what follows, we discuss these opportunities by each direction of multi-modal misinformation detection study.
6.1. Opportunities in Multi-modal Data Study
@@
Considering the data study challenges we discussed earlier, we propose the following avenues:
@@
  • Comprehensive multi-modal and multi-lingual datasets: as we discussed earlier, one important gap in the misinformation detection study is the lack of a comprehensive multi-modal dataset, which needs to be addressed in the future. Multi-modal misinformation detection requires large, multi-lingual, multi-source datasets that cover a variety of modalities, web resources, events, etc., and provide fine-grained ground truth for the samples.
@@
  • Standardized annotation strategy: Current datasets are annotated by various fact-checking agencies, leading to subjective labels in many cases. Establishing a standardized labeling agreement across all datasets would facilitate easier cross-dataset comparison and analysis.
6.2. Opportunities in Multi-modal Feature Study
@@
Based on the feature study challenges we discussed in the previous section, we propose the following research opportunities to overcome some of the existing challenges in multi-modal feature study:
  • Identifying cross-modal clues: Currently, cross-modal cues are restricted to a few basic indicators, such as the similarity between text and images. Identifying more subtle and often overlooked cues can aid in developing discordance-aware models and help recognize vulnerabilities in the serving platforms, which is integral to adversarial learning.
@@
  • Developing efficient fusion mechanisms:many of the existing solutions leverage naive fusion mechanisms such as concatenation, which may result in inefficient and noisy multi-modal representations. Therefore, another fruitful avenue of research lies in the study and development of more efficient fusion techniques to produce information-rich representations.
@@
  • Identifying language-independent features to cope with mono-lingual datasets: a majority of existing datasets are mono-lingual, thereby not sufficient enough to train models for non-English tasks. One way to compensate for the lack of multi-lingual datasets is to use language-independent features (Vogel and Meghana, 2020). Identifying such features, especially in multi-modal environments where there are more features and aspects, would be highly effective in coping with mono-lingual datasets.
6.3. Opportunities in Multi-modal Model Study
@@
Some unexplored research avenues to tackle existing model-related challenges in multi-modal misinformation detection include:
@@
  • Utilizing foundation models and prompt-based techniques in multi-modal misinformation detection: The astounding effectiveness of foundation models and techniques, including ICL and prompt-tuning, in numerous multi-modal tasks suggests that foundation models have a lot of potential for identifying multi-modal misinformation. Developing task-specific foundation models for detecting misinformation is another opportunity that would hugely impact the field of misinformation detection.
@@
  • Developing cross-modal discordance-aware architectures: most of the existing works either blindly merge modalities or take a trial-and-error approach to attend to the relevant modalities. Implementing discordance-aware models not only results in information-rich representations but also may be useful in making attention-based techniques more efficient.
@@
  • Adversarial learning in multi-modal misinformation detection: Although there are existing generative-based architectures, adversarial study of multi-modal misinformation detection has been mostly neglected. To make the detection models more adversarially robust, it is of utmost importance to dedicate time and effort to the study and development of generative and adversarial learning techniques.
@@
  • Transferable models to unseen events: As mentioned earlier, except for a few works, most of the existing models are designed for specific events and, as a result, are ineffective for emerging ones. Since misinformation spreads during a variety of events, developing general and transferable models is extremely crucial.
@@ 7. Conclusions
In this paper, we review the literature on multi-modal misinformation detection, discuss its strengths and weaknesses, and suggest new directions for future research. First, we introduce some of the prominent misinformation categories and often-used cross-modal cues for spotting them. We also discuss different fusion mechanisms to merge modalities that are engaged in such cross-modal cues. In addition, we categorize existing solutions into two groups: classic machine learning and deep learning solutions, and then further divide each group based on the techniques that are utilized. Furthermore, we introduce and compare existing datasets on multi-modal misinformation detection and identify some of the weaknesses of these datasets. By classifying them into data, feature, and model-based shortcomings, we demonstrate some of the most prominent problems in multi-modal fake news detection. Finally, we propose new lines of research to address these shortcomings.